# MAYA: Addressing Inconsistencies in Generative Password Guessing through a Unified Benchmark

### v2

@@ 4.1 Datasets Selection
TABLE I: Details of the selected datasets.
Dataset #Pass #Unique Loc Lang Year Service Rockyou [51] 32.600.024 14.311.994 USA EN 2009 Gaming Linkedin [2] 60.650.662 60.591.405 Global EN 2012 Social Mail.ru [1] 3.723.472 2.260.454 RU RU 2014 Mail 000webhost [11] 15.269.739 10.587.879 USA EN 2015 Forum Taobao [56] 7.492.035 6.165.957 CHN ZH 2012 Ecomm Gmail [41] 4.912.520 3.122.573 RU RU 2014 Mail Ashley Madison [66] 375.846 375.738 CA EN 2015 Social Libero [17] 667.680 418.400 IT IT 2016 Mail
Table I, presents the datasets selected for this study. To ensure that the selected datasets provide generalizable insights, the following criteria were considered in the selection process:
@@
We included datasets of varying sizes to evaluate model performance under different dimensions. The collection ranges from small datasets with a few hundred thousand passwords, such as Libero [17], to large-scale datasets with tens of millions of passwords, such as RockYou [51], LinkedIn [2] and 000webhost [11].
@@
The selected datasets vary in their geographical provenance. Previous studies have shown that users from different countries exhibit distinct password creation behaviors, leading to diverse password distributions [58, 36, 18, 23, 59].
@@
We selected datasets representing diverse languages and cultural contexts, as users from different backgrounds exhibit variations in password choices [58, 5, 36, 59].
@@
We selected datasets from a wide range of services and communities, such as social networks, forums, e-commerce sites, dating sites, email services, and gaming platforms. The type of service and the nature of the user community can significantly influence password selection strategies, with more sensitive service and close-knit communities often prompting users to adopt different password behaviors [54, 3, 30].
@@ 4.2 Dataset Analysis
TABLE II: Distribution of password lengths.
@@ 4.2.1 Password Length Analysis
Table II shows that the most common password lengths are between 6 and 10 characters. Taobao, 000webhost, LinkedIn, and Libero enforce a stricter policy (len >= 6), with very few passwords shorter than 6 characters. The Cumulative Distribution Function (CDF) reveals that over $50\%$ of passwords are at most 8 characters long, with 8 being, on average, the most frequently chosen length. Notably, the CDF rapidly increases up to 12 characters, suggesting that passwords longer than 12 characters are rare. Based on these observations, we selected passwords with maximum lengths of 8, 10, and 12 in our experiments.
@@ 4.2.2 Password Patterns Analysis
Based on common password rules and user behaviors, we defined 19 patterns, listed in Table III, to characterize password patterns across the datasets. For a detailed distribution analysis, see Table X in the Appendix. In most datasets, letter-only passwords are highly prevalent and are almost always exclusively lowercase, whereas uppercase-only or mixed-case passwords appear infrequently, apart from 000webhost, likely influenced by a stricter password policy enforced prior to the breach. In European and American datasets, digit-only passwords are less frequent than letter-only ones, whereas Taobao (from China) shows a higher prevalence of digit-only passwords, consistent with prior studies [58, 5, 36]. This is likely due to users being less familiar with the Latin alphabet. Passwords comprised entirely of special characters are rarely used, as they are hard to remember, and users prioritize usability over security. A common pattern across all datasets is the combination of letters and digits. In 000webhost $\approx 93\%$ of passwords follow this pattern, suggesting a policy requiring at least two character classes. Users often find password policies frustrating [53, 52, 32] and tend to fall into predictable patterns to comply with these requirements, such as appending a trailing digit [58, 59]. Our analysis supports this observation, as passwords ending with a digit are highly prevalent, with nearly half of all passwords following this pattern and ‘1’ being widely used as a final character.
TABLE III: Selected patterns and their description.
@@
Figure 1: Password frequency distribution. The x-axis value is computed over the total number of passwords without duplicates.
@@ 4.2.3 Top-10 Passwords Analysis
We analyzed the top 10 passwords in each dataset, finding that common choices like ‘123456’ appear across almost all datasets. Notably, password preferences vary by region: European users favor lowercase-only passwords, Chinese users often choose digit-based passwords resembling words phonetically, and Russian users prefer keyboard patterns. Table XI in the Appendix presents the top 10 most common passwords.
@@ 5.1 RQ1 - How Do Different Settings Influence Models Performance?
We examine the impact of two key parameters—maximum password length and the number of generated passwords—on model performance. Specifically, we compute the weighted average percentage of guessed passwords across all datasets, considering maximum lengths of 8, 10, and 12 characters. For each length, we evaluate 11 generation quantities, ranging from $10^{6}$ to $5\times 10^{8}$ passwords. Results are shown in Figure 2. As expected, performance declines with increasing password length. However, this decline exhibits diminishing returns: while the drop from 8 to 10 characters is substantial, averaging 10.46%, the decrease from 10 to 12 characters is markedly smaller, at just 2.15%, indicating a plateau as length increases. Among all models, FLA consistently outperforms the others, successfully guessing a substantial portion of passwords with relatively few guesses. Notably, FLA requires fewer than $10^{7}$ generated passwords to outperform all other models except PassGPT. PassGPT, while initially comparable to PLR-GAN, PassFlow, and PassGAN, exhibits a much steeper growth curve, ultimately closing the gap with FLA after $5\times 10^{7}$ guesses. When considering few guesses, PLR-GAN, PassFlow, and PassGAN exhibit nearly identical performance, while VGPT2 initially underperforms. As the number of guesses increases, differences between the models become more pronounced: PLR-GAN begins to outperform PassFlow, and VGPT2, despite a slow start, improves faster than both. In contrast, PassGAN gains the least from additional guesses, with a relatively slow improvement rate.
While it is evident that generating more passwords leads to an increased number of matches, we delve deeper into this trend by analyzing the marginal gain—defined as the percentage increase in guessed passwords between two generation intervals. We analyze both total gains (relative to the overall number of correct matches) and relative gains (relative to the previous number of matches). Due to space constraints, detailed results are reported in Table XII in the Appendix. Overall, all models exhibit clear diminishing returns, with sub-linear growth in successful guesses. Match rates grow rapidly in the early stages, then taper off as the space of common passwords is exhausted. Since this pattern is consistent across all maximum lengths, and considering that over $92\%$ of passwords are 12 characters or fewer (see Table II), we limit our subsequent analysis to this length.
@@
TABLE IV: Impact of training dataset size on performance. Values expressed as percentage of guessed test set passwords.
Train Size Mailru Taobao RockYou LinkedIn PGAN PLR PFLW PGPT VGPT FLA PGAN PLR PFLW PGPT VGPT FLA PGAN PLR PFLW PGPT VGPT FLA PGAN PLR PFLW PGPT VGPT FLA 1e6 12.73 21.26 23.27 32.93 6.25 53.14 13.15 16.45 17.58 27.51 4.24 42.94 9.60 19.29 18.21 30.98 5.97 56.13 4.42 7.83 7.06 12.49 1.48 18.48 2e6 16.16 21.30 23.24 40.35 16.90 54.95 10.16 16.29 18.65 29.11 11.28 44.12 11.31 19.21 18.35 41.72 13.75 57.36 4.38 8.08 7.12 17.97 4.44 24.61 3e6 10.80 16.05 17.21 29.81 11.37 44.79 12.78 17.69 19.32 44.01 15.25 60.45 3.79 7.92 6.88 19.75 4.55 29.30 5e6 12.11 16.82 18.49 30.80 12.56 45.53 12.22 18.06 19.35 46.40 16.49 60.08 2.52 8.38 6.84 22.00 5.38 37.89 1e7 12.36 19.65 18.46 48.85 17.90 60.47 3.33 6.53 7.00 25.27 5.97 41.05 2e7 4.87 7.12 7.13 26.91 6.33 35.10 4e7 3.99 8.44 7.10 28.58 6.56 36.37
5.2 RQ2 - Are Generative Models Truly Better Than Traditional Tools?
We compare generative models against two categories of traditional password-guessing techniques: (1) rule-based attacks, namely, Hashcat and John the Ripper, both operating in wordlist mode, using ’Unicorn Rules’ and ’Wordlist’ rulesets, respectively; and (2) machine learning approaches, specifically a PCFG model and OMEN, a Markov-based attack. Additional details on these tools are provided in Appendix A.
@@ 5.2 RQ2 - Are Generative Models Truly Better Than Traditional Tools?
However, as dataset size increases, the advantage shifts toward generative and machine-learning-based models, with the performance gap over rule-based tools widening considerably. On average, Hashcat underperforms compared to the top four learning-based approaches, while JtR consistently ranks as the least effective. We further investigated the rationale behind this declining performance, hypothesizing it stems from their underlying generation strategies. Unicorn Rules are ordered by efficacy, and we selected the first $n$ passwords generated by Hashcat accordingly. However, Hashcat applies a rule-first strategy: it processes all rules for the current password before moving to the next. As the dataset size increases, a smaller portion of the test set passwords are transformed, concentrating Hashcat’s guesses in a small subset of the overall password space. JtR follows a password-first strategy, but WordList’s rules are unsorted, so we randomly sampled $n$ passwords from the generated set. In larger datasets, the number of generated candidates increases substantially, diluting the proportion of successful guesses and making it less likely to sample correct guesses.
@@
TABLE V: Cross-community generalization ability. Values expressed as percentage of guessed test set passwords.
Train / Test PassGAN PLR-GAN PassFlow PassGPT VGPT2 FLA Link Rock  [^0] W Link Rock  [^0] W Link Rock  [^0] W Link Rock  [^0] W Link Rock  [^0] W Link Rock  [^0] W Linkedin 3.99 6.51 1.80 8.44 12.65 3.65 7.10 6.61 1.93 28.58 36.21 16.90 6.56 11.48 2.95 36.37 45.09 19.01 RockYou 4.72 12.36 1.59 8.77 19.65 4.65 8.16 18.46 3.55 22.08 48.85 13.15 6.93 17.90 2.84 31.53 60.47 17.31  [^0] Webhost 2.27 4.10 2.63 3.70 5.56 5.22 6.58 11.44 2.69 10.41 13.77 20.24 1.87 3.15 2.59 22.81 26.72 32.28
TABLE VI: Cross-culture generalization ability. Values expressed as percentage of guessed test set passwords.
Train / Test PassGAN PLR-GAN PassFlow PassGPT VGPT2 FLA Rock. Mail. Taob. Rock. Mail. Taob. Rock. Mail. Taob. Rock. Mail. Taob. Rock. Mail. Taob. Rock. Mail. Taob. RockYou 12.36 8.43 5.57 19.65 13.22 9.10 18.46 14.86 9.83 48.85 22.30 13.65 17.90 11.40 6.52 60.47 30.10 18.71 Mailru 7.39 16.16 4.72 10.78 21.30 7.11 16.35 23.24 13.25 15.82 40.35 10.74 6.89 16.90 4.11 26.98 54.95 16.36 Taobao 6.41 7.39 12.11 9.37 10.11 16.82 19.55 20.20 18.49 13.26 13.61 30.80 7.74 8.16 12.56 23.77 20.94 45.53

### v4

@@ 4.1 Datasets Selection
TABLE II: Details of the selected datasets.
General Filtering Policy Dataset #Password #Unique Loc Lang Year Service #Rem %Rem Char Len Rockyou [59] 32.600.024 14.311.994 USA EN 2009 Gaming 17.788 0.05% NONE $\geq$ 5 Linkedin [2] 60.650.662 60.591.405 Global EN 2012 Social 0 0.00% NONE $\geq$ 6 Mail.ru [1] 3.723.472 2.260.454 RU RU 2014 Mail 0 0.00% NONE $\geq$ 5 000webhost [13] 15.269.739 10.587.879 USA EN 2015 Forum 18.742 0.12% 2 Classes $\geq$ 6 Taobao [64] 7.492.035 6.165.957 CHN ZH 2012 Ecomm 5 0.00% NONE $\geq$ 6 Gmail [49] 4.912.520 3.122.573 RU RU 2014 Mail 10.453 0.21% NONE - AshleyMad. [75] 375.846 375.738 CA EN 2015 Social 0 0.00% NONE - Libero [20] 667.680 418.400 IT IT 2016 Mail 331 0.05% NONE $\geq$ 6
Table II presents the datasets selected for this study. For ethical reasons, we avoided newer and larger datasets sold in unverified forums or websites and focused on publicly available ones used in previous works. To ensure that the selected datasets provide generalizable insights, the following criteria were considered in the selection process:
@@
We included datasets of varying sizes to evaluate model performance under different dimensions. The collection ranges from small datasets with a few hundred thousand passwords, such as Libero [20], to large-scale datasets with tens of millions of passwords, such as RockYou [59], LinkedIn [2], and 000webhost [13]. We excluded archives comprising aggregations of multiple datasets, as they would introduce biases in generalizability analysis and may overlap with the other datasets used.
@@
The selected datasets vary in their geographical provenance. Previous studies have shown that users from different countries exhibit distinct password creation behaviors, leading to diverse password distributions [66, 43, 21, 26, 67]. This offers an opportunity to assess models across different password patterns.
@@
We selected datasets representing diverse languages and specific cultural contexts. This allows us to study the generalizability between different cultures, as users from different backgrounds exhibit variations in password choices [66, 5, 43, 67].
@@
We selected datasets from a wide range of services and communities, such as social networks, forums, e-commerce sites, dating sites, email services, and gaming platforms. While some datasets are highly community-specific, they were specifically selected to assess generalizability across niche or very different communities. The type of service and the nature of the user community can significantly influence password selection strategies, with more sensitive services and close-knit communities often prompting users to adopt different password behaviors [62, 3, 35].
@@ 4.2 Dataset Analysis
TABLE III: Distribution of password lengths.
@@ 4.2.1 Password Length Analysis
Table III shows that the most common lengths are between 6 and 10 characters. Taobao, 000webhost, LinkedIn, and Libero enforce a stricter policy (len >= 6), with very few passwords shorter than 6 characters. The Cumulative Distribution Function (CDF) reveals that over $50\%$ of passwords are at most 8 characters long, with 8 being, on average, the most frequently chosen length. Notably, the CDF rapidly increases up to 12 characters, suggesting that passwords longer than 12 characters are rare. Based on these observations, we selected passwords with maximum lengths of 8, 10, and 12 in our experiments.
@@ 4.2.2 Password Patterns Analysis
Based on common password rules and user behaviors, we defined 19 patterns, listed in Table IV, to characterize password patterns across the datasets. For a detailed distribution analysis, see Table XIII in the Appendix. In most datasets, letter-only passwords are highly prevalent and are almost always exclusively lowercase, whereas uppercase-only or mixed-case passwords appear infrequently, apart from 000webhost, likely influenced by a stricter password policy enforced prior to the breach. In European and American datasets, digit-only passwords are less frequent than letter-only ones, whereas Taobao (from China) shows a higher prevalence of digit-only passwords, consistent with prior studies [66, 5, 43]. This is likely due to users being less familiar with the Latin alphabet. Passwords comprised entirely of special characters are rarely used, as they are hard to remember, and users prioritize usability over security. A common pattern across all datasets is the combination of letters and digits. In 000webhost $\approx 93\%$ of passwords follow this pattern, suggesting a policy requiring at least two character classes. Users often find password policies frustrating [61, 60, 39] and tend to fall into predictable patterns to comply with these requirements, such as appending a trailing digit [66, 67]. Our analysis supports this observation, as passwords ending with a digit are highly prevalent, with nearly half of all passwords following this pattern and ‘1’ being widely used as a final character.
TABLE IV: Selected patterns and their description.
@@
Figure 1: Password frequency distribution. The x-axis is computed over the total number of unique passwords.
@@ 4.2.3 Top-10 Passwords Analysis
We analyzed the top 10 passwords in each dataset, finding that common choices like ‘123456’ appear across almost all datasets. Notably, password preferences vary by region: European users favor lowercase-only passwords, Chinese users often choose digit-based passwords resembling words phonetically, and Russian users prefer keyboard patterns. Table XIV in the Appendix presents the top 10 most common passwords.
@@ 4.2.4 Frequency Distribution
TABLE V: Marginal gain in successful matches as the number of generated passwords increases. Each column represents a generation interval (e.g., 1M-2.5M indicates the marginal gain from 1M to 2.5M guesses). Total reports the marginal gain from X to Y, expressed as a percentage of the total number of test passwords. Relative indicates the marginal gain relative to the number of matches obtained at X. Results are averaged across all password lengths.
Model 1M-2.5M 2.5M-5M 5M-7.5M 7.5M-10M 10M-25M 25M-50M 50M-75M 75M-100M 100M-250M 250M-500M Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative FLA +4.33 +58.54 +4.17 +35.55 +2.66 +16.71 +1.93 +10.40 +6.57 +32.14 +5.55 +20.62 +3.33 +10.28 +2.34 +6.56 +7.44 +19.62 +5.84 +13.00 OMEN +1.50 +66.31 +1.20 +30.96 +1.20 +23.46 +1.56 +25.29 +2.46 +32.46 +4.02 +37.08 +2.94 +20.95 +1.17 +6.94 +7.25 +38.16 +6.05 +22.75 PassFlow +0.62 +69.67 +0.73 +48.88 +0.56 +25.15 +0.44 +15.72 +1.72 +52.72 +1.58 +31.53 +1.05 +15.85 +0.80 +10.38 +2.84 +33.66 +2.48 +21.97 PassGAN +0.49 +83.75 +0.55 +50.94 +0.40 +24.84 +0.32 +15.95 +1.23 +53.74 +1.14 +32.85 +0.75 +16.47 +0.56 +10.69 +1.96 +34.15 +1.63 +21.38 PassGPT +1.20 +114.72 +1.60 +71.91 +1.34 +35.32 +1.18 +22.97 +5.21 +83.58 +5.45 +48.49 +3.66 +22.15 +2.73 +13.56 +8.89 +39.13 +6.53 +20.63 PCFG +2.08 +37.87 +3.19 +42.00 +2.85 +26.96 +1.88 +14.08 +6.27 +41.03 +4.30 +19.99 +2.40 +9.33 +1.68 +5.95 +5.09 +17.06 +3.58 +10.25 PLR-GAN +0.62 +97.50 +0.74 +59.45 +0.57 +28.94 +0.48 +18.91 +1.94 +65.41 +1.96 +40.34 +1.34 +19.54 +1.03 +12.55 +3.67 +40.28 +3.21 +25.59 VGPT2 +0.18 +137.88 +0.28 +90.08 +0.25 +43.08 +0.24 +28.34 +1.18 +110.51 +1.45 +65.38 +1.11 +30.52 +0.92 +19.44 +3.65 +65.02 +3.43 +37.03
@@ 5.1 RQ1 - How Do Different Settings Influence Models Performance?
We examine the impact of two key parameters—maximum password length and the number of generated passwords—on model performance. Specifically, we compute the weighted average percentage of guessed passwords across all datasets, considering maximum lengths of 8, 10, and 12 characters. For each length, we evaluate 11 generation quantities, ranging from $10^{6}$ to $5\times 10^{8}$ passwords. Results are shown in Figure 2. As expected, performance declines with increasing password length. However, this decline exhibits diminishing returns: while the drop from 8 to 10 characters is substantial, averaging 10.71 percentage points, the decrease from 10 to 12 characters is markedly smaller, at just 2.80 points, indicating a plateau as length increases. Among all models, FLA consistently outperforms the others, successfully guessing a substantial portion of passwords with relatively few guesses. Notably, FLA requires fewer than $10^{7}$ generated passwords to outperform all other models except PassGPT, PCFG, and OMEN. PassGPT, while initially comparable to PLR-GAN, PassFlow, and PassGAN, exhibits a much steeper growth curve, growing even faster than FLA after $5\times 10^{7}$ passwords. When considering few guesses, PLR-GAN, PassFlow, and PassGAN exhibit nearly identical performance, while VGPT2 initially underperforms. As the number of guesses increases, differences between the models become more pronounced: PLR-GAN begins to outperform PassFlow, and VGPT2, despite a slow start, improves faster than both. In contrast, PassGAN gains the least from additional guesses, with a relatively slow improvement rate. OMEN and PCFG, perform surprisingly well. OMEN initially follows a curve similar to PassGPT but later falls behind, especially for longer passwords. PCFG’s curve initially follows FLA’s, but its performance gains diminish as more guesses are generated. At $5\times 10^{8}$ guesses, PCFG performs worse than PassGPT and OMEN for length 8, and slightly better than PassGPT for lengths 10 and 12.
While it is evident that generating more passwords leads to an increased number of matches, we delve deeper into this trend by analyzing the marginal gain—defined as the percentage increase in guessed passwords between two generation intervals. We analyze both total gains (relative to the overall number of correct matches) and relative gains (relative to the previous number of matches). Detailed results are reported in Table V. Overall, all models exhibit clear diminishing returns, with sub-linear growth in successful guesses. Match rates grow rapidly in the early stages, then taper off as the space of common passwords is exhausted. Since this pattern is consistent across all lengths, and considering that over $92\%$ of passwords are 12 characters or fewer (see Table III), we limit our subsequent analysis to this length.
@@
TABLE VI: Impact of training dataset size on performance. Values expressed as percentage of guessed test set passwords.
Size Mailru Taobao RockYou LinkedIn FLA OMEN PFLW PGAN PGPT PCFG PLR VGPT FLA OMEN PFLW PGAN PGPT PCFG PLR VGPT FLA OMEN PFLW PGAN PGPT PCFG PLR VGPT FLA OMEN PFLW PGAN PGPT PCFG PLR VGPT 1e6 53.14 37.78 23.27 12.73 32.93 36.24 21.26 6.25 42.94 27.90 17.58 13.15 27.51 19.10 16.45 4.24 56.13 38.70 18.21 9.60 30.98 32.80 19.29 5.97 18.48 12.68 7.06 4.42 12.49 22.89 7.83 1.48 2e6 54.95 38.42 23.24 16.16 40.35 40.26 21.30 16.90 44.12 28.18 18.65 10.16 29.11 22.05 16.29 11.28 57.36 39.65 18.35 11.31 41.72 36.67 19.21 13.75 24.61 13.07 7.12 4.38 17.97 25.33 8.08 4.44 3e6 44.79 28.19 17.21 10.80 29.81 23.75 16.05 11.37 60.45 39.93 19.32 12.78 44.01 38.71 17.69 15.25 29.30 13.22 6.88 3.79 19.75 26.45 7.92 4.55 5e6 45.53 28.29 18.49 12.11 30.80 26.17 16.82 12.56 60.08 40.11 19.35 12.22 46.40 41.27 18.06 16.49 37.89 13.32 6.84 2.52 22.00 27.80 8.38 5.38 1e7 60.47 40.29 18.46 12.36 48.85 45.24 19.65 17.90 41.05 13.37 7.00 3.33 25.27 30.37 6.53 5.97 2e7 35.10 13.42 7.13 4.87 26.91 31.94 7.12 6.33 4e7 36.37 13.42 7.10 3.99 28.58 33.69 8.44 6.56
5.2 RQ2 - Are Deep Generative Models Truly Better Than Traditional Tools?
We compare DL-based generative models against two categories of traditional password-guessing techniques: (1) rule-based attacks, namely, Hashcat and John the Ripper, both operating in wordlist mode, using ’Unicorn Rules’ and ’Wordlist’ rulesets, respectively; and (2) ML-based generative approaches, specifically PCFG and OMEN. Additional details on these tools are provided in Appendix A.
@@ 5.2 RQ2 - Are Deep Generative Models Truly Better Than Traditional Tools?
However, as dataset size increases, the advantage shifts toward learning-based models, with the performance gap over rule-based tools widening considerably. On average, Hashcat underperforms compared to the top four learning-based approaches, while JtR consistently ranks as the least effective. We further investigated the rationale behind this declining performance, hypothesizing it stems from their underlying generation strategies. Unicorn Rules are ordered by efficacy, and we selected the first $n$ passwords generated by Hashcat accordingly. However, Hashcat applies a rule-first strategy: it processes all rules for the current password before moving to the next. As the dataset size increases, a smaller portion of the test set passwords are transformed, concentrating Hashcat’s guesses in a small subset of the overall password space. JtR follows a password-first strategy, but WordList’s rules are unsorted, so we randomly sampled $n$ passwords from the generated set. In larger datasets, the number of generated candidates increases substantially, diluting the proportion of successful guesses and making it less likely to sample correct guesses.
@@
TABLE VII: Cross-community generalization ability. Values expressed as percentage of guessed test set passwords.
Train / Test FLA OMEN PassFlow PassGAN PassGPT PCFG PLR-GAN VGPT2  [^0] W. Link. Rock.  [^0] W. Link. Rock.  [^0] W. Link. Rock.  [^0] W. Link. Rock.  [^0] W. Link. Rock.  [^0] W. Link. Rock.  [^0] W. Link. Rock.  [^0] W. Link. Rock.  [^0] Webhost 31.33 22.81 26.72 11.72 7.55 9.63 2.74 6.58 11.44 2.66 2.27 4.10 20.24 10.41 13.77 28.78 20.30 23.93 5.22 3.70 5.56 2.59 1.87 3.15 LinkedIn 19.01 36.37 45.09 7.53 13.42 17.90 1.93 7.10 6.61 1.80 4.03 6.51 16.90 28.58 36.21 23.99 33.69 40.30 3.65 8.44 8.77 2.95 6.56 11.48 RockYou 17.31 31.53 60.47 8.10 17.60 40.29 3.55 8.16 18.46 1.59 4.72 12.41 13.15 22.08 48.85 20.75 27.17 45.24 4.65 8.77 19.67 2.84 6.93 17.90
TABLE VIII: Cross-culture generalization ability. Values expressed as percentage of guessed test set passwords.
Train / Test FLA OMEN PassFlow PassGAN PassGPT PCFG PLR-GAN VGPT2 Mail. Rock. Taob. Mail. Rock. Taob. Mail. Rock. Taob. Mail. Rock. Taob. Mail. Rock. Taob. Mail. Rock. Taob. Mail. Rock. Taob. Mail. Rock. Taob. Mailru 54.95 26.98 16.36 38.42 15.89 12.29 23.37 16.35 13.25 16.29 7.39 4.72 40.35 15.82 10.74 40.26 14.57 6.37 21.32 10.78 7.11 16.90 6.89 4.11 RockYou 30.10 60.47 18.71 19.43 40.29 16.28 14.86 18.48 9.83 8.43 12.41 5.57 22.30 48.85 13.65 23.25 45.24 9.54 13.22 19.67 9.10 11.40 17.90 6.52 Taobao 20.94 23.77 45.53 11.11 10.05 28.29 20.20 19.55 18.68 7.39 6.41 12.16 13.61 13.26 30.80 10.64 11.72 26.17 10.11 9.37 16.84 8.16 7.74 12.56